OpenStudy (anonymous):

How do you factor ax^2+bx+c

OpenStudy (anonymous):

Can you please give specific numbers? The following only shows how quadratic equation is like.

OpenStudy (anonymous):

\[a(x-r_1)(x-r_2)\] were \[r_1=\frac{-b+\sqrt{b^2-4ac}}{2a}\] and \[r_2=\frac{-b-\sqrt{b^2-4ac}}{2a}\]